# Do Robots Really Need Anthropomorphic Hands?

### v1

@@ 3.3.1 Planning
Motion planning [43] and grasp planning, utilizing tools such as GraspIt! [60] and grasp quality measures [81], are traditional approaches for generating robot motion. However, they often assume a noise-free, controlled environment, and it is challenging to define arbitrary manipulation tasks beyond grasping or pick-and-place.
3.3.2 Dynamical Movement Primitives
An approach to robot control that combines aspects of generalized motor programs and schema theory on the one hand and dynamical systems theory on the other is dynamical movement primitives (DMPs, [85]). DMPs define stable dynamical systems to generate motor commands, and can be learned from demonstration and through reinforcement learning. Similar to the concept of schema theory, regression techniques can be used to parameterize DMPs based on context or task parameters.
3.3.3 Stochastic Optimal Feedback Control, Optimal Control, Model-Predictive Control, and Reinforcement Learning
@@ 3.3.2 Dynamical Movement Primitives
Optimal control has been implemented in robotic systems to generate movements in various forms. [65] generate complex manipulation actions for various hands through optimization with an a priori defined model of the system and environment. However, a similar approach can be used continuously online with a high frequency, which is known as model-predictive control, to generate complex movements as shown by [99] in the context of locomotion. This approach is robust to model errors.
In model-based RL, dynamics models and reward models are learned from experience. [67] demonstrate how to use a model for online optimization of actions in various manipulation tasks, e.g., in-hand manipulation. Model-free RL solves the problem without explicitly learning the dynamics model. The most successful approaches for continuous control [e.g., 73, 120] are based on the actor-critic framework, which directly learns a policy that maps observations to actions, eliminating the need for online optimization. RL can solve complex tasks, such as playing piano [120].
Although these approaches are promising and attempt to mimic human motor control, in practice, they often fail to integrate complex sensory data. They could, but even though camera images are sometimes used directly as input, they are often avoided [e.g., 67] or enriched with additional position information [e.g., 73]. In comparison to human motor control, we often overlook the integration of additional sensors, such as tactile sensors. In addition, learning and adaptation of behavior typically do not occur online, even with machine learning approaches, after the initial training stage is completed.
3.3.4 Vision-Language-Action Models
Motivated by the success of large pretrained vision and language models, vision-language-action models (e.g., RT-2 [122]) trained on teleoperation data are a novel imitation learning approach for visuo-motor skills. Although this learning paradigm is still in its early stage of research, it appears promising, as integrating other sensor modalities seems possible.
@@ 3.4.2 Sensors
From a perceptual perspective, human hands have a vast array of sensors, including mechanoreceptors, thermoreceptors, and nociceptors, which provide detailed information about touch, texture, temperature, and potential for damage (Section 2.2). This sensory input is essential to perform a wide range of object manipulations, allowing humans to adjust their grip forces and hand or finger positions in real time. In contrast, robotic hands rely on a limited number of tactile or force sensors, which are often bulky compared to the size of the robotic fingers and provide only partial information about the environment. Additionally, the data from these sensors requires post-processing to be interpreted and used effectively in manipulation tasks. These limitations suggest the need for further development in both the design and signal processing aspects of robotic perception sensors.
@@ 3.4.3 Control and Learning
We have seen rapid progress in robot control recently, with groundbreaking works on control of complex hands [e.g., 73]. However, the integration of complex sensory feedback and complex hands remains an open challenge, and online learning and adaptation are not the primary focus. Although the control approaches are similar to those used by humans, they do not reach human dexterity.
@@ 4 Systematic Review of Robotic Hands and their Skill Repertoire
We performed a systematic literature review to connect hands with the skills they can perform. We use a simplified version of the PRISMA approach [61] as we did not aggregate the results of empirical studies with reported statistical results, but merely collect information on demonstrated robot skills.
@@ 4.1.2 Selection
We filtered the results from Google Scholar and IEEE Xplore using the following criteria: publication dates must be between 2019 and 2025, the article must have at least one citation, be at least four pages long, and be written in the English language. All IEEE papers were removed from the initial set retrieved using Google Scholar to eliminate duplicates. All non-peer-reviewed articles published on arXiv were removed from our list. Finally, we filtered IEEE papers based on the keyword terms to ensure they contain one of these terms: (hand or gripper) and (manipulation or grasping or grip or skill).
@@ 4.1.3 Data collection process
We examined the papers and extracted any skills or movements that the authors claim to have generated for a real or simulated robotic hand. These manipulation skills were then categorized according to the taxonomy of manipulation tasks by [21]. The taxonomy encompasses both prehensile skills, such as grasping and in-hand manipulation, and non-prehensile skills, including pushing and gestures. It is designed to be independent of the hand morphology. Furthermore, for in-hand manipulation categories (14 and 15: Contact / Motion / Within Hand / (No) Motion at Contact), we extract the degrees of freedom of an object that are manipulated.
@@ 4.1.4 Risk of bias in individual studies
We see the risk of underreported skill repertoires for the robotic hands. Not all hands fully exhaust their skill repertoire in the respective publication. Even though we aggregate results from various publications using the same hand, we can not be sure that the hand would not be able to perform an in-hand manipulation skill with more DoF or another category of skills not seen in the publication. Thus, it is likely that some skills that a hand would be able to perform are not investigated in the respective publication. For instance, the users of prosthetic hands may be able to perform a large variety of skills in their daily lives that are not reported in any publication. As a result, we analyze extreme cases, e.g., the minimum complexity of a hand required to implement a specific skill.
@@ 4.2.2 Synthesis of results
Outliers that we identified in Figure 15 are (1) [12] with 2 fingers, 2 DoF, 1 actuator, and 3 controllable DoF of the object, as well as (2) the following examples of 6 controllable DoF of the object: two hands of [66] with 3 fingers, 7 DoF, 4 actuators, and 4 fingers, 8 DoF, 4 actuators respectively, and [55] with 4 fingers, 8 DoF, 4 actuators.
Figure 15: Number of actuators versus degrees of freedom that could be controlled in one of the in-hand manipulation categories. Extreme cases: (1) 3 controllable DoF of the object: [12] with one actuator (2 fingers, 2 DoF); (2) 6 controllable DoF of the object: two hands of [66] with four actuators (3 fingers, 7 DoF; 4 fingers, 8 DoF), and [55] with four actuators (4 fingers, 8 DoF).
@@ 4.3.1 Summary of evidence
We could not find a correlation between hand complexity and the DoFs of an object that the hand is able to manipulate.
@@
We focused on the outliers [12, 66, 55] and found that they can control many DoF of an object without complex mechanisms.
@@ 4.3.3 Conclusions
The overall complexity of integrated hand mechanisms and sensors does not yet reach the level of human capabilities. Specifically, the integration of many sensors and complex mechanisms is underexplored. It is technically feasible to assemble a Shadow Dexterous Hand with tactile sensors and use a camera, but it is challenging to control intelligently.
The entire skill repertoire of most hands remains underexplored, and we recognize a need for more thorough and standardized evaluation of both new hands and existing hands.
There is no correlation between hand complexity and skill complexity for probably two reasons:
(1) The most complex hands in the analyzed studies (5 fingers, >20>20 degrees of freedom, and >12>12 actuators) were often capable of more complex in-hand manipulation skills than demonstrated. However, most studies use these hands to manipulate 1–3 degrees of freedom only. Just four exceptions manipulate $\geq 4$ DoF of an object.
(2) Simple, non-anthropomorphic mechanisms are sufficient to manipulate more degrees of freedom than we expected, as demonstrated by the identified outliers [12, 66, 55]. When analyzing these more closely, we see that [12] use the simplest possible gripper design with two fingers, two DoF, and one actuator. They use the environment intelligently as an additional contact point to manipulate the object, which could be interpreted as using an additional virtual finger. [66] and [55] use non-anthropomorphic designs. Their hands have four fingers arranged in a circular shape around a center point, so that an object can be controlled from four different directions and enabling control of six DoF during in-hand manipulation.
@@ 5.1 Mechanical Features of the Human Hand
Many tasks can be solved by simple mechanisms that have two fingers and one actively controlled degree of freedom. The gripper of [12] that can perform in-hand manipulation with 3 DoF by using the environment as a third contact surface in a clever way. [54] build upon this idea. Similarly, humans constantly exploit environment constraints [20]. For this to be possible, robotic hands must be robust, i.e., flexible and soft, with complexity shifting to intelligent motor control and sensor integration. [95] argue that animals that lack complex hands are still capable of complex manipulation. For instance, birds can use their bills to grasp, regrasp, and adjust the grasp of objects to establish a stable grip and gather information about the object, as well as exhibit complex behavior, such as building nests. [63] observe in experiments that many human grasps can be performed with the thumb and a single virtual finger. Additional DoFs in the human hand were only necessary for precision grasps, which are used less often in activities of daily living.
@@
An additional finger is sometimes needed. Parrots have a tongue with which they can rotate objects in their bills or stabilize them during manipulation [95], essentially realizing 1 DoF in-hand manipulation with three finger-like body parts. Similarly, [121] found that the three-fingered BarrettHand can perform functional grasps to use tools such as an electric drill or a flashlight. [1] provide further evidence that three fingers are sufficient for most tasks. They use the grasp taxonomy of [24] to analyze the force distribution among fingers and palm in human grasps and found that only three fingers exert force in most cases. Other fingers are only used to fine-tune the grasp. Furthermore, [25] show that two or three fingers have an increased workspace volume and rotational range when compared to more fingers for the specific case of in-hand precision manipulation of circular objects.
Even in domains where we would expect multi-fingered hands to excel, the evidence is not clear. We present examples from deformable object manipulation and multi-object grasping.
@@
Deformable objects constantly deform under external forces, e.g., the forces applied by a robotic hand. The more actively controllable contact points or surfaces a gripper has with a deformable object, the more control it has over the object’s shape. Interestingly, [83] found that much research on deformable objects has been conducted with two fingers and they suggest that using more fingers allows for finer control. This motivated [22] to develop a gripper for the tasks of picking up and placing folded clothes, folding a t-shirt in the air, and tracing an edge of a grasped cloth. They found that using three fingers with one finger opposing two others is particularly useful for the task of picking and placing folded clothes as abduction of the two fingers at the bottom creates a large support surface for the clothes that would otherwise collapse around a thin support surface. Hence, even in this domain, in which we assume that more fingers and DoFs are better, three fingers may be sufficient.
@@
[96] distinguish twelve multi-object grasp types from human demonstrations and a robot grasping data set, in which eight different object types were grasped. They replicated many of these grasps using a BarrettHand with three fingers, which demonstrates that three fingers can be sufficient for complex grasps involving multiple objects. However, in some cases, this was not possible, e.g., the inverse basket grasp, in which five fingers form a basket around the objects and apply grip force, and the multi-finger pinch grasp, in which multiple pinch grasps were used simultaneously on multiple objects.
@@
These observations suggest that a wide variety of manipulation tasks, including simple in-hand manipulation, can be solved with two fingers or at most three fingers for additional grasp stabilization during manipulation or tool use. However, [95] note that birds make use of a highly flexible neck that functions as an equivalent to the human wrist during manipulation, and [62] make the equivalent observation in human manipulation with restricted DoFs in their hand. Robotic wrists are often not as flexible, and orientation changes often require large repositioning of the arm, which makes operation in restricted workspaces difficult.
@@
[66] and [55] use a non-anthropomorphic design with two pairs of opposing fingers so that 6 DoF of an object can be controlled during in-hand manipulation. This mechanism can be equally dexterous as an anthropomorphic hand and might even be better suited for specific in-hand manipulation tasks.
@@
Humans use two hands for many tasks. An example of a complex manipulation task is tying shoe laces, which requires two pairs of opposing fingers. [59] report that individuals with six fully functional fingers can solve this task with only one hand. As the research of [96] on multi-object grasps suggests, more fingers allow handling multiple objects simultaneously and conforming better with the shape of multiple objects. We conclude that the more tasks that have to be solved in parallel, the more fingers are required. Complex in-hand manipulation can benefit from more fingers, and human in-hand manipulation is likely limited not just by the degrees of freedom per finger and range of motion per joint, but also by the number of fingers, as has been demonstrated by [59].
@@
How many fingers should we use? Almost certainly, robots do not need five fingers like the human hand. Currently, there is no evidence that five fingers are optimal. This raises questions about which configuration is most effective. Two fingers can be sufficient for many tasks. Moreover, the range of tasks can even be extended with the effective use of environment contacts. Three fingers are a good compromise between dexterity and simplicity, as they improve stability and enable tool use. A flexible wrist significantly enhances manipulation capabilities considerably for two and three fingers. Four fingers provide more dexterity for in-hand manipulation, and with two more fingers, we found evidence of increasing dexterity. If and how more fingers can be used to improve robotic manipulation remains an open question.

### v2

@@ 3.3.1 Planning
Traditional approaches for generating robot motion are motion planning [93] and grasp planning, utilizing tools such as GraspIt! [94] and grasp quality measures [95]. However, these often assume a noise-free, controlled environment, and it is challenging to define arbitrary manipulation tasks beyond grasping or pick-and-place.
3.3.2 Based on Motor Program Theory and Dynamical Systems Theory
An approach to robot control that combines aspects of generalized motor programs and schema theory on the one hand and dynamical systems theory on the other is dynamical movement primitives (DMPs, [96]). DMPs define stable dynamical systems to generate motor commands, and can be learned from demonstration in imitation learning and through reinforcement learning. Similar to the concept of schema theory, regression techniques can be used to parameterize DMPs based on context or task parameters.
The idea of muscle synergies has been adopted in robotic control. For instance, Ben Amor et al. [97] extract a low-dimensional grasp space with principal component analysis from human demonstrations to control a multi-fingered robotic hand for reaching and grasping.
3.3.3 Based on Stochastic Optimal Feedback Control
@@
Optimal control has been implemented in robotic systems to generate movements in various forms. Mordatch et al. [98] generate complex manipulation actions for various hands through optimization with an a priori defined model of the system and environment. However, a similar approach can be used continuously online with a high frequency, which is known as model-predictive control, to generate complex movements as shown by Tassa et al. [99] in the context of locomotion. This approach is robust to model errors.
In model-based RL, dynamics models and reward models are learned from experience. Nagabandi et al. [100] demonstrate how to use a model for online optimization of actions in various manipulation tasks, e.g., in-hand manipulation. Model-free RL solves the problem without explicitly learning the dynamics model. The most successful approaches for continuous control [e.g., 59, 101] are based on the actor-critic framework, which directly learns a policy that maps observations to actions, eliminating the need for online optimization. RL can solve complex tasks, such as playing piano [101].
Although these approaches are promising and attempt to mimic human motor control, in practice, they often fail to integrate complex sensory data. They could, but even though camera images are sometimes used directly as input, they are often avoided, e.g., due to the curse of dimensionality or partial observability [e.g., 100] or enriched with additional position information [e.g., 59]. In comparison to human motor control, we often overlook the integration of additional sensors, such as tactile sensors. In addition, learning and behavioral adaptation typically do not occur online after the initial training stage is completed, even with machine learning approaches.
3.3.4 Imitation Learning for Sensorimotor Skills
Currently, a more promising category of algorithms to obtain visuo-motor skills is imitation learning, in which control policies are learned from expert demonstrations. For instance, Chi et al. [102] demonstrate this for tabletop manipulation with various simple end-effectors and grippers. Zhao et al. [103] extend this to challenging bimanual manipulation tasks with a simple two-finger gripper: replacing a robot finger, hanging a shirt, and tying shoelaces. Success rates for these tasks vary between 40% and 75%, depending on the difficulty.
Motivated by the success of large pretrained vision-language models (VLMs), vision-language-action models [104] are a novel imitation learning approach for visuo-motor skills. Although this learning paradigm is still in its early stage of research, it appears promising, as integrating other sensor modalities and more complex hands is possible. Unlike previous imitation learning approaches, these models aim to be generalists, solving various tasks as well as complex, high-level tasks that are composed of multiple subtasks. Similarly, large behavior models [105], which are trained on large-scale robot data, but not necessarily based on a pretrained VLM, are a promising future research direction that could produce generalist sensorimotor control.
@@ 3.4.2 Sensors
From a perceptual perspective, human hands have a vast array of sensors, including mechanoreceptors, thermoreceptors, and nociceptors, which provide detailed information about touch, texture, temperature, and potential for damage (Section 2.2). This sensory input is essential to perform a wide range of object manipulations, allowing humans to adjust their grip forces and hand or finger positions in real time. In contrast, robotic hands rely on a limited number of tactile or force sensors, which are often bulky relative to the robotic fingers and provide only partial information about the environment. Additionally, the data from these sensors requires post-processing to be interpreted and used effectively in manipulation tasks. These limitations suggest the need for further development in at least the mechanical and electrical design and signal-processing aspects of robotic perception sensors.
@@ 3.4.3 Control and Learning
We have seen rapid progress in robot control recently, with groundbreaking research on control of complex hands [e.g., 59]. However, integrating complex sensory feedback and complex hands remains an open challenge, and online learning and adaptation are not the primary focus yet. Although the control approaches are similar to those used by humans, they do not reach human dexterity.
3.4.4 Prosthetic Hands
@@ 4 Systematic Review of Robotic Hands and their Skill Repertoire
We performed a systematic literature review to connect hands with the skills they can perform. We use a simplified version of the PRISMA approach [107] because we did not aggregate the results of empirical studies with reported statistical results. Instead, we limited ourselves to collecting information on demonstrated robot skills due to the lack of standardization of design goals or requirements, and, more importantly, testing protocols.
@@ 4.1.2 Selection
We filtered the results from Google Scholar and IEEE Xplore using the following criteria: publication dates must be between 2019 and 2025, the article must have at least one citation, be at least four pages long, and be written in English. All IEEE papers were removed from the initial set retrieved using Google Scholar to eliminate duplicates. All non-peer-reviewed articles published on arXiv were removed from our list. Finally, we filtered IEEE papers based on the keyword terms to ensure they contained one of the following terms: (hand or gripper) and (manipulation or grasping or grip or skill).
@@ 4.1.3 Data collection process
We examined the papers and extracted any skills or movements that the authors claim to have generated for a real or simulated robotic hand. These manipulation skills were then categorized according to the taxonomy of manipulation tasks by Dollar [108]. The taxonomy encompasses both prehensile skills, such as grasping and in-hand manipulation, and non-prehensile skills, including pushing and gestures. It is designed to be independent of the hand morphology. Furthermore, for in-hand manipulation categories (14 and 15: Contact / Motion / Within Hand / (No) Motion at Contact), we extract the degrees of freedom of an object that are manipulated.
@@ 4.1.4 Risk of bias in individual studies
We see the risk of underreported skill repertoires for the robotic hands. Not all hands fully exhaust their skill set in their respective publication. Even though we aggregate results from various publications using the same hand, we cannot be sure that the hand would not be able to perform an in-hand manipulation skill with more DoF or another category of skills not seen in the publications. Thus, it is likely that some skills a hand can perform are not investigated in the respective publication. For instance, users of prosthetic hands may be able to perform a large variety of skills in their daily lives that are not reported in any publications. As a result, we analyze extreme cases, e.g., the minimum complexity of a hand required to implement a specific skill. Another limitation is the binary nature of reporting skills, which means it is not possible to infer the quality or proficiency of performing a skill with a particular hand.
@@ 4.2.2 Synthesis of results
Outliers that we identified in Figure 13 are (1) Chavan-Dafle et al. [109] with 2 fingers, 2 DoF, 1 actuator, and 3 controllable DoF of the object, as well as (2) the following examples of 6 controllable DoF of the object: two hands of Morgan et al. [110] with 3 fingers, 7 DoF, 4 actuators, and 4 fingers, 8 DoF, 4 actuators respectively, and Liarokapis and Dollar [111] with 4 fingers, 8 DoF, 4 actuators.
Figure 13: Number of actuators versus degrees of freedom that could be controlled in one of the in-hand manipulation categories. Extreme cases: (1) 3 controllable DoF of the object: Chavan-Dafle et al. [109] with one actuator (2 fingers, 2 DoF); (2) 6 controllable DoF of the object: two hands of Morgan et al. [110] with four actuators (3 fingers, 7 DoF; 4 fingers, 8 DoF), and Liarokapis and Dollar [111] with four actuators (4 fingers, 8 DoF).
@@ 4.3.1 Summary of evidence
We could not find a correlation between hand complexity and the DoFs of an object that the hand is able to manipulate. Reasons for this are that simple mechanisms are sufficient to control many DoFs and most complex hands do not fully exhaust their capabilities.
@@
We focused on the outliers [109, 110, 111] and found that they can control many DoF of an object without complex mechanisms by using the environment as additional virtual finger [109] or using two pairs of opposing fingers [110, 111].
@@ 4.3.3 Conclusions
The overall complexity of integrated hand mechanisms and sensors does not yet reach the level of human capabilities. Specifically, the integration of many sensors and complex mechanisms is underexplored. It is technically feasible to assemble a Shadow Dexterous Hand with tactile sensors and use a camera, but it is still challenging to control intelligently.
The entire skill repertoire of most hands remains underexplored, and we recognize a need for more thorough and standardized evaluation of both new hands and existing hands. There is no correlation between hand complexity and skill complexity for probably two reasons: (1) The most complex hands in the analyzed studies (5 fingers, >20>20 degrees of freedom, and >12>12 actuators) were often capable of more complex in-hand manipulation skills than demonstrated. However, most studies use these hands to manipulate 1–3 degrees of freedom only. Just four exceptions manipulate $\geq 4$ DoF of an object. (2) Simple, non-anthropomorphic mechanisms are sufficient to manipulate more degrees of freedom than we expected, as demonstrated by the identified outliers [109, 110, 111]. When analyzing these more closely, we see that Chavan-Dafle et al. [109] use the simplest possible gripper design with two fingers, two DoF, and one actuator. They use the environment intelligently as an additional contact point to manipulate the object, which could be interpreted as using an additional virtual finger. Morgan et al. [110] and Liarokapis and Dollar [111] use non-anthropomorphic designs. Their hands have four fingers arranged in a circular shape around a center point, so that an object can be controlled from four different directions and enabling control of six DoF during in-hand manipulation.
@@ 5.1 Mechanical Features of the Human Hand
Many tasks can be solved by simple mechanisms that have two fingers and one actively controlled degree of freedom. The gripper of Chavan-Dafle et al. [109] that can perform in-hand manipulation with 3 DoF by using the environment as a third contact surface in a clever way. Liang et al. [112] build upon this idea. Similarly, humans constantly exploit environment constraints [113]. For this to be possible, robotic hands must be robust, i.e., flexible and soft, with complexity shifting to intelligent motor control and sensor integration. Sugasawa et al. [2] argue that animals that lack complex hands are still capable of complex manipulation. For instance, birds can use their bills to grasp, regrasp, and adjust the grasp of objects to establish a stable grip and gather information about the object, as well as exhibit complex behavior, such as building nests. While their grippers often only have two virtual fingers, their neck is flexible. Montagnani et al. [114] observe in experiments that many human grasps can be performed with the thumb and a single virtual finger. Additional DoFs in the human hand were only necessary for precision grasps, which are used less often in activities of daily living. Zhao et al. [103] show that complex tasks, such as tying shoe laces, can be solved with two two-fingered robotic grippers. Although their success rates vary between 40% and 70% for this task, this may be because of the difficulty of generating an appropriate visuomotor control policy and not because of the simplicity of the gripper.
@@
A third finger improves functional grasping and in-hand manipulation performance. Zhu et al. [115, 116] found that the three-fingered BarrettHand can perform functional grasps to use tools such as an electric drill or a flashlight. Similarly, Parrots have a tongue with which they can rotate objects in their bills or stabilize them during manipulation [2], essentially realizing 1 DoF in-hand manipulation with three virtual fingers. In particular, Zhu et al. [115] compare functional grasps of the five-fingered Shadow dexterous hand and the three-fingered Barrett hand and demonstrate that the fingers of each hand correctly touch the functional area of the objects (e.g., screwdriver, flashlight, pincers, stapler) while maintaining a stable grasp. Abbasi et al. [117] provide further evidence that three fingers are sufficient for most tasks. They use the grasp taxonomy of Feix et al. [118] to analyze the force distribution among fingers and palm in human grasps and found that only three fingers exert force in most cases. Other fingers are only used to fine-tune the grasp. This underlines that the additional finger considerably improves manipulation abilities and increasing the number of fingers further is not necessarily required.
Feix et al. [119] further show that two or three fingers in contact with an object have an increased workspace volume and rotational range when compared to more fingers in contact for the specific case of in-hand precision manipulation of circular objects with the human hand. They demonstrate that with an increasing number of digits in contact with the object, the manipulation workspace range for four DOF of the object reduces significantly. They point out that two digits result in the largest workspace in five of the six DOF, however, three digits are required to control the missing rotational DOF. Hence, they argue that three fingers result in the best overall performance for in-hand manipulation.
@@
Even in domains where we would expect multi-fingered hands to excel, the evidence is unclear. We present examples from deformable object manipulation and multi-object grasping. Deformable objects change their shape under external forces, e.g., those applied by a robotic hand. The more actively controllable contact points or surfaces a gripper has with a deformable object, the greater its control over the object’s shape. Still, Sanchez et al. [120] find that considerable research on deformable objects was conducted with two fingers and they argue that using more fingers allows for finer control. This motivated Donaire et al. [121] to develop a gripper for picking up and placing folded clothes, folding a t-shirt in the air, and tracing the edge of a grasped cloth. They found that using three fingers, with one finger opposing two others, is particularly useful for picking and placing folded clothes, as abduction of the two bottom fingers creates a large support surface for the clothes that would otherwise collapse around a thin support surface. Hence, even in this domain, where we assume that more fingers and DoFs are better, three fingers already provide a considerable improvement over two fingers.
@@
Sun et al. [122] distinguish twelve multi-object grasp types from human demonstrations and a robot grasping data set, in which eight different object types were grasped. They replicated many of these grasps using a BarrettHand with three fingers, which demonstrates that three fingers can be sufficient for complex grasps involving multiple objects. However, in some cases, this was not possible, e.g., the inverse basket grasp, in which five fingers form a basket around the objects and apply grip force, and the multi-finger pinch grasp, in which multiple pinch grasps were used on multiple objects. While the inverse basket grasp is designed for five fingers, it is not clear that five individually controllable fingers have an advantage over two virtual fingers with larger surface. Only in the case of the multi-finger pinch grasp, the hand obviously benefits from more fingers, as they enable grasping more objects simultaneously.
@@
These observations suggest that a wide variety of manipulation tasks, including simple in-hand manipulation, can be solved with two fingers or at most three fingers for additional grasp stabilization during manipulation or tool use. However, Sugasawa et al. [2] note that birds make use of a highly flexible neck that functions as an equivalent to the human wrist during manipulation, and Montagnani et al. [123] make the equivalent observation in human manipulation with restricted DoFs in their hand.
@@
Morgan et al. [110] and Liarokapis and Dollar [111] use a non-anthropomorphic design with two pairs of opposing fingers so that 6 DoF of an object can be controlled during in-hand manipulation. This mechanism can be equally dexterous as an anthropomorphic hand and might even be better suited for specific in-hand manipulation tasks.
@@
Humans use two hands for many tasks. An example of a complex manipulation task is tying shoe laces, which requires two pairs of opposing fingers. Mehring et al. [3] report that individuals with six fully functional fingers can solve this task with only one hand. As the research of Sun et al. [122] on multi-object grasps suggests, more fingers allow handling multiple objects simultaneously and conforming better with the shape of multiple objects. We conclude that the more tasks that have to be solved in parallel, the more fingers are required. Complex in-hand manipulation can benefit from more fingers, and human in-hand manipulation is likely limited not just by the degrees of freedom per finger and range of motion per joint, but also by the number of fingers, as has been demonstrated by Mehring et al. [3].
Fingers and wrist flexibility
The human arm possesses seven DOF: three in the shoulder, one in the elbow, and three in the wrist. Hence, we have a redundant DOF (see Section 3.1), which allows us to incorporate secondary objectives besides pose control, e.g., avoiding obstacles or torque limits [55, 56]. Major et al. [124] find that lack of wrist flexibility is compensated by arm and trunk movements in human subjects. Furthermore, Montagnani et al. [123] show that wrist flexibility is more important than finger dexterity in current prosthetics. In their experiments with constraint human hands, they find that restricting the DOFs of the human wrist to pronation/supination (removing two DOFs) results in similar performance on the SHAP benchmark as restricting the hand function to open/close (removing 21 DOF). Similarly, Sugasawa et al. [2] stress that flexible necks are essential for birds’ manipulation capabilities.
@@
How many fingers should we use? Almost certainly, robots do not need five fingers like the human hand. Currently, there is no evidence that five fingers are optimal. This raises questions about which configuration is most effective. Two fingers can be sufficient for many tasks. Moreover, the range of tasks can even be extended with the effective use of environment contacts. Three fingers often are a good compromise between dexterity and simplicity, as they improve stability and enable tool use. A flexible wrist enhances manipulation capabilities considerably for two and three fingers. Four fingers provide more dexterity for in-hand manipulation, and with two more fingers, we found evidence of increased dexterity over five fingers for tasks that usually be solved bimanually. How more than five fingers can be used effectively to improve robotic manipulation remains an open research question.